OpenStudy (anonymous):

Suzie worked the following problem incorrectly. In which line was the mistake made and what is the correct answer? Line 1 50 / 5(3 − 1) − 4 Line 2 50 / 5(2) − 4 Line 3 50 / 10 − 4 Line 4 5 - 4 Line 5 1 :(

OpenStudy (anonymous):

can someone explain how to do this...

OpenStudy (anonymous):

WOAH!

OpenStudy (anonymous):

Is the -4 on the right supposed to be in the denominator or is it a separate term?

OpenStudy (anonymous):

LINE 2

OpenStudy (anonymous):

wait no...

OpenStudy (anonymous):

goodman im suppose to write the question out and im not sure how to explain it

OpenStudy (anonymous):

i dunt understand line 3 and 4..so the problem must be in that group...

OpenStudy (anonymous):

\[\frac{50}{5(3-1)}-4\] \[\frac{50}{5\times 2}-4\] \[\frac{50}{10}-4\] \[5-4\] \[1\] suzie aced this one for sure!

OpenStudy (anonymous):

but it said she was incorrect...

OpenStudy (anonymous):

The only thing I can see wrong with that is if the -4 should be in the denominator. If that's the case, the error is in line 3.

OpenStudy (anonymous):

yeah she said it was incorrect

OpenStudy (anonymous):

ya..i agreee with coderiferous and mike...

OpenStudy (anonymous):

ok what line is wrong and how would you fix th wrong line???

OpenStudy (anonymous):

like coderiferous said..its in line3

OpenStudy (anonymous):

line 3 is wrong so how would you explain how yo correct it?

OpenStudy (anonymous):

Yeah, but only if the -4 is part of the denominator, which I have to assume it is.

OpenStudy (anonymous):

hmm this is a hard one to do

OpenStudy (anonymous):

It would be 50/(10-4) which is 50/6, or 25/3

OpenStudy (anonymous):

so ''It would be 50/(10-4) which is 50/6, or 25/3'' whould be explaing it right?

OpenStudy (anonymous):

Can you use the equation function to recreate the exact lines, or are the lines written out exactly how you typed?

OpenStudy (anonymous):

there exactly like that nothing is changed